a) Given the demand function for computers to be P = 2400 – 0.4Q


i. Determine the coefficient of point elasticity of demand when P = 1800, P = 1200,


and P = 600 and give an interpretation of each result. [7]


ii. If the price of computers increases by 12%, calculate the percentage change in


quantity demanded at P = 1800, P = 1200, and P = 600. [12]

1
Expert's answer
2021-11-30T02:40:30-0500

i.


"P=2400 \u2013 0.4Q"

"1=-0.4\dfrac{dQ}{dP}"

"\dfrac{dQ}{dP}=-2.5"

"\eta=\dfrac{dQ}{dP}\times\dfrac{P}{Q}"

"\eta=-2.5\times\dfrac{P}{Q}"

"P=1800, 1800=2400 \u2013 0.4Q=>Q=1500"


"\eta=-2.5\times\dfrac{1800}{1500}=-3"

Therefore, at this point on the demand curve, a 1 percent change in price causes a 3 percent change in quantity demanded in the opposite direction (because of the negative sign).


"P=1200, 1200=2400 \u2013 0.4Q=>Q=3000"


"\eta=-2.5\times\dfrac{1200}{3000}=-1"

Therefore, at this point on the demand curve, a 1 percent change in price causes a 1 percent change in quantity demanded in the opposite direction (because of the negative sign).


"P=600, 600=2400 \u2013 0.4Q=>Q=4500"

"\eta=-2.5\times\dfrac{600}{4500}=-1/3"

Therefore, at this point on the demand curve, a 1 percent change in price causes a 1/3 percent change in quantity demanded in the opposite direction (because of the negative sign).


ii.


"P=1800, 12\times(-3)=-36"

If the price of computers increases by 12%, the percentage change in quantity demanded decreases by 36%.



"P=1200, 12\times(-1)=-12"

If the price of computers increases by 12%, the percentage change in quantity demanded decreases by 12%.



"P=600, 12\times(-1/3)=-4"

If the price of computers increases by 12%, the percentage change in quantity demanded decreases by 4%.